Web Service

web62, web63, web64, web65 abnormal servers performance

Last Updated
15 August 2017 - 07:17 (UTC)

Aug 15, 08:17 BST
Resolved - This incident has been resolved.

Aug 15, 07:50 BST
Investigating - We are currently having adverse performance on these servers, and are investigating the causes.
We expect to have them back to normal operation soon, and apologise for any inconvenience caused.

Exchange Platform Connectivity

Last Updated
14 August 2017 - 10:22 (UTC)

Aug 14, 11:22 BST
Resolved - This incident has been resolved.

Aug 14, 11:04 BST
Identified - We have identified a connectivity issue with our hosted Exchange platform. We are currently working to resolve this issue as soon as possible. We apologize for any inconvenience caused by this.

Mail114 abnormal server performance

Last Updated
12 August 2017 - 12:20 (UTC)

Aug 12, 13:20 BST
Resolved - This incident has been resolved.

Aug 12, 11:26 BST
Identified - We are currently having adverse performance on this server which is affecting the processing and delivery of emails coming into this mail-server. We are investigating the cause and expect to have it back to normal operation as soon as possible. We apologise for any inconvenience caused.

web135, web136, web137, web138 abnormal server behaviour

Last Updated
12 August 2017 - 12:20 (UTC)

Aug 12, 13:20 BST
Resolved - This incident has been resolved.

Aug 12, 08:24 BST
Update - Update - The restore is currently at 99.80% complete

Aug 11, 22:55 BST
Update - The restore is currently at 86.16% complete

Aug 11, 21:31 BST
Update - The restore is currently at 81.98% complete

Aug 11, 20:06 BST
Update - The restore is currently at 77.49% complete

Aug 11, 17:49 BST
Update - The restore is currently at 71.70% complete

Aug 11, 15:41 BST
Update - The restore is currently at 68.90% complete.

Aug 11, 14:31 BST
Update - The restore is currently at 67.40% complete.

Aug 11, 13:36 BST
Update - The restore is currently at 65.89% complete.

Initially we expected the restoration process to take around 24 hours to be complete, but we are now seeing an increase in the estimated time until completion, as the sites being restored are incrementally larger in size.

Our current estimate for the restore is 1 AM tomorrow. We will update you as soon as we have more information.

Aug 11, 12:45 BST
Update - The restore is currently at 64.70% complete.

Aug 11, 11:21 BST
Update - The restore is currently at 61.68% complete.

Aug 11, 09:44 BST
Update - The restore is currently at 58.38% complete.

Aug 11, 08:14 BST
Update - The restore is currently at 54.81% complete.

Aug 11, 07:01 BST
Update - The restore is currently at 51.67% complete.

Aug 11, 06:01 BST
Update - The restore is currently at 47.62% complete.

Aug 11, 05:03 BST
Update - The restore is currently at 45.45% complete.

Aug 11, 04:08 BST
Update - The restore is currently at 43.43% complete.

Aug 11, 03:01 BST
Update - The restore is currently at 41.40% complete.

Aug 11, 01:45 BST
Update - The restore is currently at 39.58% complete.

Aug 11, 00:49 BST
Update - The restore is currently at 37.34% complete.

Aug 10, 23:24 BST
Update - The restore is currently at 34.85% complete.

Aug 10, 22:35 BST
Update - The restore is currently at 34.05% complete.

Aug 10, 21:37 BST
Update - The restores are currently on 31%, the ETA remains unchanged currently.

Aug 10, 19:19 BST
Update - The restorations are currently at 23% complete, we will update on this progress throughout the night.

Aug 10, 18:01 BST
Update - 18% of all the packages on this cluster have been restored now. These sites are back online and should be fully functional.

We still expect this to finish according to the ETA provided. We'll update you as soon as we have more information.

Aug 10, 17:06 BST
Update - The restoration process is 15% complete now and all sites that have been restored so far are back online. There are still no changes to the ETA provided for this restoration.

We'll update you as soon as we have more information.

Aug 10, 16:07 BST
Update - The restoration is still progressing as planned on the NAS drives, and there are no changes to the ETA at the moment.

We are prioritising small sized sites at the moment in order to ensure that the majority of sites on these servers are back online as quickly as possible.

We'll update you as soon as we have more information.

Aug 10, 14:19 BST
Identified - An issue has been identified with the NAS drives for the above servers. All sites on these servers are effected by this issue, but the databases remain intact.

We are restoring the NAS drives from our backups and we expect this restoration process to take around 24 hours.

We'll update you as soon as we have more information, we apologise for any inconvenience caused regarding this.

Aug 10, 11:42 BST
Investigating - We are currently seeing abnormal behaviour on this server, and are investigating the causes.
We expect to have it back to normal operation soon, and apologise for any inconvenience caused.

Close